# How Do Document Parsers Break? Auditing Structural Vulnerability in Document Intelligence

### 19 May 2026 v1

@@ 5.1 Experimental Setup
We instantiate $\mathcal{F}=(\mathcal{S},\Pi,\mathcal{D})$ on 1,000 validation pages from PubLayNet (Zhong et al., 2019) and DocLayNet (Pfitzmann et al., 2022), after filtering pages with fewer than five original spans.
@@
The main text reports configuration-level analyses, with image fixed-effects regression, per-image OLS, Spearman analysis, dose-response binning, and within-configuration quantile checks reported in Appendix C.6.
Setting MinerU $P$-V3
Input images / valid pages 1,000 1,000
Shared evaluation pool 1,000
Complexity (S / M / C) 14.5% / 46.4% / 39.1%
Phase 1 configs / Phase 2 policies 29 / 5
Matching thresholds $\tau_{\mathrm{iou}}\!=\!0.1,\ \tau_{\mathrm{text}}\!=\!0.5$
Table 1: Dataset and protocol summary.
@@ 5.2.2 Downstream Impact
Table 2: Downstream propagation. All values are percentages; EM Drop and AnsMiss are lower-better degradation signals, while EM and AnsHit@5 are higher-better task scores; AM/Str./LA denote area-matched/structural/large-area; bold marks the more degraded AM–Str. result.
@@
Table 2 reports a controlled propagation test for downstream document use. Rather than introducing a separate document-QA benchmark, we reuse the same 975 clean-derived QA pairs under four counterfactual parser outputs, so that only the perturbation condition changes. For retrieval, the corpus is restricted to chunks from the same page and uses the same block-aware chunking across all conditions; this isolates whether answer-bearing evidence remains available after parsing. The key contrast controls for visual footprint: area-matched erasure (AM) and the structural probe (Str.) perturb nearly the same page area, with TOR around 0.10–0.11%.
@@ 5.2.3 B-SLR Channels and Faithfulness
Returning to the parser-level diagnostic, we next examine whether composite B-SLR is internally decomposable. Under the joint IoU+text criterion, each failed element is assigned to either an IoU-fail channel, where no valid geometric match survives, or a text-fail-only channel, where geometry is retained but textual identity is lost. The two channels sum to composite B-SLR, and Table 3 reports their configuration-level means and coupling statistics over the 29 Phase 1 configurations.
The decomposition reveals two patterns. First, text-side failure is non-redundant and dominant in both pipelines: its mean contribution exceeds the IoU channel on MinerU (.043 vs. .028) and more strongly on PP-V3 (.085 vs. .026). Thus, many structural failures preserve coarse geometry while losing recognized content. Second, the channel coupling differs sharply across architectures. MinerU shows coherent geometry-text variation ($\rho_{s}=.891$), whereas PP-V3 is driven mainly by the text channel ($R^{2}_{\mathrm{text}}=.900$ vs. $R^{2}_{\mathrm{IoU}}=.186$). This suggests that PP-V3 concentrates vulnerability in recognition-side failure, while MinerU distributes failure across both geometry and text.
These channel-level differences also clarify how B-SLR should be interpreted. Because CER measures perturbation-induced text instability, the text channel is expected to be informative; the useful observation is that the geometry–text coupling differs sharply by architecture rather than collapsing to a single text-only score. Composite B-SLR therefore serves as an output-level structural diagnostic aligned with the terminal OCR-instability signal, reaching $R^{2}=.727/.916$ and Spearman $\rho=.911/.973$ on MinerU/PP-V3.
5.2.4 Gradient Predicts Failure Pathways
@@
Parser Policy Exposure / Budget Terminal & Structural Damage Failure Pathway Efficiency
@@
Table 5: Phase 2 policy audit. $\Delta\mathrm{mAP}=\mathrm{mAP}(\mathcal{Y},E)-\mathrm{mAP}(\mathcal{Y},E_{\mathrm{adv}})$; $\mathrm{TopoShare}=\mathrm{SLR}_{\mathrm{topo}}/(\mathrm{SLR}_{\mathrm{miss}}+\mathrm{SLR}_{\mathrm{topo}})$; $\mathrm{Eff}_{\mathrm{B\mbox{-}SLR}}=\mathrm{B\mbox{-}SLR}/\mathrm{TOR}$, $\mathrm{Eff}_{\mathrm{CER}}=\mathrm{CER}/\mathrm{TOR}$. $\uparrow$: denotes larger per-footprint efficiency
@@ 5.2.5 Sensitivity Across Probe Families
The NT sweep further reveals parser-dependent sensitivity to sustained structural interference. MinerU increases gradually as the target interference level $\tau$ grows, whereas PP-StructureV3 shows a steeper rise in topology-dominant loss at higher target levels. Overall, Phase 1 shows that the probe space contains qualitatively different failure regimes: severe erasure produces miss-heavy spikes, while structure-targeted interference increasingly induces topology-dominant failure.
@@ 5.2.6 Targeting Efficiency Across Policies
Table 5 reports the Phase 2 policy-driven audit, where all policies share the same probe space $\mathcal{S}$ and page pool. We use these policies not as competing attacks, but as probes of whether targeting logic can expose structural damage under different area budgets.
The results reinforce the footprint-bias finding under policy-level selection. Structurally targeted policies often reach comparable structural loss with much smaller area budgets: for example, Rule-based matches Random on PP-V3 B-SLR despite using about one seventh of the TOR, and exceeds Random on MinerU CER with a similarly small budget. Thus, policy choice changes not only how much area is perturbed, but whether the selected area is structurally consequential.
The pathway decomposition shows that this efficiency is not mainly direct content removal. TopoShare remains high (.80–1.00), and $\mathrm{SLR}_{\mathrm{miss}}$ is near zero for most non-random policies. Absolute damage and per-footprint efficiency therefore expose complementary aspects of vulnerability: Random can produce larger raw damage by spending more area, whereas targeted policies reveal structure-sensitive failure per unit footprint.
@@ 6 Conclusion
We introduced a lightweight, output-level tripartite auditing framework for structure-aware DLA vulnerability auditing. Across MinerU and PP-StructureV3, B-SLR tracks OCR instability more faithfully than footprint, exposure descriptors separate occlusion- and topology-dominant pathways, and targeted probes reveal disproportionate downstream QA/retrieval degradation under small area budgets. Overall, DLA reliability should be evaluated structurally, not by footprint alone.
@@
This work focuses on diagnostic auditing rather than universal robustness certification. First, our experiments cover two representative DLA pipelines with different detection and recognition designs. This setting is sufficient to expose cross-architecture differences, but broader coverage across additional parsers, detector families, and end-to-end document models would further test the generality of the observed diagnostic patterns. Second, our framework requires block-level boxes and recognized text, so systems that expose only free-form textual outputs or latent structured representations require an additional alignment interface before they can be audited in the same way. Third, our mechanistic interpretation is based on controlled black-box interventions and output-level diagnostics. While this is appropriate for pipeline-level auditing, white-box analyses of detector and recognizer internals would be needed to localize the exact feature-level causes of failure. Fourth, per-pixel efficiency metrics such as B-SLR/TOR are most informative when comparing probes within comparable area ranges; they should be interpreted together with absolute damage and pathway composition. Finally, although the evaluation pool combines PubLayNet and DocLayNet pages, dataset-specific annotation conventions may affect absolute scores, and larger cross-domain audits remain an important extension.
@@ A.1 Probe Space and Probe Catalog
Probe Geom. Behavior Key parameters Placement $\mathcal{T}$
P1 (horiz. crease) line inject $w\!\in\![1,10],\;l_{r}\!\in\![0.5,1.0]$ a/c/r
P2 (vert. crease) line inject $w\!\in\![1,10],\;l_{r}\!\in\![0.5,1.0]$ a/c/r
P4 (rect. modification) rect erase $a\!\in\![3\%,25\%],\;\beta\!\in\![0.2,1.0]$ c/b
P5 (thin horiz. line) line inject $w\!\in\![1,5],\;l_{r}\!\in\![0.2,0.8]$ b/c/r
P8 (irregular patch) blob blend $r_{b}\!\in\![30,80],\;\epsilon\!\in\![0.1,0.5],\;\alpha\!\in\![0.3,0.7]$ a
@@ A.2 Probe Generation and Placement
This subsection specifies how the probe catalog in Table 6 is rendered on a document page. We denote geometry masks by $\mathcal{H}_{(\cdot)}$ to distinguish probe geometry from the diagnostic gradient families in the main text. The geometry primitives are implemented as binary masks:
$\mathcal{H}_{\mathrm{line}}(\theta,l,w)=\left\{(x,y):|x\sin\theta-y\cos\theta|<\tfrac{w}{2},\right$.
$\qquad\left.0\leq x\cos\theta+y\sin\theta\leq l\right\},$ (21)
@@
$\mathcal{H}_{\mathrm{rect}}(w_{r},h_{r})=\{(x,y):|x|\leq w_{r}/2,\;|y|\leq h_{r}/2\},$ (23)
$\mathcal{H}_{\mathrm{blob}}(r_{b},\epsilon)=\left\{(r,\phi):r\leq r_{b}\bigl(1+\epsilon\cdot\mathrm{Perlin}(\phi)\bigr)\right\},$ (24)
$\mathcal{H}_{\mathrm{point}}(n,r,\sigma)=\bigcup_{i=1}^{n}\mathcal{H}_{\mathrm{disk}}(r)\big|_{\text{center}\sim\mathcal{N}(\mu,\sigma^{2}I)}.$ (25)
@@
$I^{\prime}(x,y)=\bigl(1-\alpha(x,y)\bigr)I(x,y)+\alpha(x,y)t(x,y),$ (26)
where $t(x,y)$ is the probe texture or background target. The three composition behaviors in Table 6 are instantiated as inject ($\alpha=1$), blend ($\alpha\in(0,1)$), and erase (blending toward the local background color).
@@
$M_{\mathrm{anchor}}=\mathcal{D}\!\left(\bigcup_{i}\partial b_{i},\delta\right)\setminus M_{\mathrm{content}}^{\mathrm{erode}(\delta)},$ (27)
@@ A.3 Policy Families
$\pi:\;c\mapsto\hat{P}=\langle\hat{\mathcal{H}},\hat{\mathcal{V}},\hat{\mathcal{B}},\hat{\mathcal{T}}\rangle$, (28)
where $c$ denotes context features extracted from the current page, and $\hat{P}$ follows the same geometry–appearance–behavior–placement schema as Section A.1.
@@ B.1 Exposure Descriptor Definitions
This subsection gives the operational definitions of the five exposure descriptors used in Section 4.4.2. Let $S(P)$ denote the perturbation mask induced by probe configuration $P$, and let $(W,H)$ be the page size. For annotated layout elements $L$, we define the annotated layout support and boundary support as
@@
$M_{\partial L}$ $=\mathcal{D}\!\left(\bigcup_{\ell\in L}\partial b(\ell),\delta\right)$
$\quad\setminus\Omega_{L}^{\mathrm{erode}(\delta)}.$
@@
$\mathrm{TOR}(P)$ $=\frac{|S(P)|}{W\times H},$ (30)
$\mathrm{ACR}(P,L)$ $=\frac{|S(P)\cap\Omega_{L}|}{|\Omega_{L}|},$
$\mathrm{BPO}(P,L)$ $=\frac{|S(P)\cap M_{\partial L}|}{|M_{\partial L}|},$
$\mathrm{BOC}(P,L)$ $=\frac{|\{\ell\in L:b(\ell)\cap S(P)\neq\emptyset\}|}{|L|},$
$\mathrm{EIR}(P,E)$ $=\frac{|\{e\in E:b(e)\cap S(P)\neq\emptyset\}|}{|E|}.$
@@ B.2 Output Matching and TextSim
$\mathrm{IoU}(e,m(e))\geq\tau_{\mathrm{iou}},\qquad\mathrm{TextSim}(e,m(e))\geq\tau_{\mathrm{text}}.$ (32)
@@
Text consistency is computed by character-level longest common subsequence:
$\mathrm{TextSim}(a,b)=\frac{|\mathrm{LCS}(a,b)|}{\max(|a|,|b|)}.$ (33)
Before computing TextSim, we strip leading/trailing whitespace and case-fold to lowercase; internal whitespace is preserved, and no punctuation removal or tokenization is applied. TextSim is used only as a matching criterion for B-SLR, not as the CER computation itself.
@@ B.4 Terminal Metric Definitions
For the OCR channel, we use character error rate (CER) as the primary terminal judge. Because the audit compares clean and perturbed parser outputs, the clean parser text serves as the reference text for each matched element. Using the maximum-IoU correspondence defined in Appendix B.2, let $r_{e}=\operatorname{txt}(e)$ and $h_{e}=\operatorname{txt}(m(e))$. The element-wise CER is
@@
$E_{\mathrm{text}}=\left\{e\in E\;\middle|\;\operatorname{txt}(e)\neq\epsilon\right\}$ (36)
@@
$\Delta_{\mathrm{det}}:=\Delta\mathrm{mAP}=\mathrm{mAP}(\mathcal{Y},E)-\mathrm{mAP}(\mathcal{Y},E_{\mathrm{adv}}).$ (39)
@@ B.5 Operational Constants and Threshold Rationale
$\rho(e,S(P))=\frac{|b(e)\cap S(P)|}{|b(e)|}$ (40)
measures the fraction of an element box directly covered by the perturbation mask. A structurally failed element with $\rho(e,S(P))\geq\eta_{\mathrm{occ}}$ is assigned to the direct-occlusion subset $\mathcal{U}_{\mathrm{miss}}$; otherwise, it is assigned to the topology-dominant subset $\mathcal{U}_{\mathrm{topo}}$.

### 26 May 2026 v2

@@ 5.1 Experimental Setup
We instantiate $\mathcal{F}=(\mathcal{S},\Pi,\mathcal{D})$ on a shared pool of 1,000 validation pages from PubLayNet (Zhong et al., 2019) and DocLayNet (Pfitzmann et al., 2022), after filtering pages with fewer than five original spans; the simple/medium/complex split is 14.5%/46.4%/39.1%.
@@
The main text reports configuration-level analyses, with image fixed-effects regression, per-image Ordinary Least Squares (OLS) regression, Spearman analysis, dose-response binning, and within-configuration quantile checks reported in Appendix C.6.
@@ 5.2.2 Downstream Impact
Table 1: Downstream propagation. All values are percentages. EM: exact match; A@5: answer hit at top 5. AM/Str./LA: area-matched/structural/large-area. Bold marks the worse AM–Str. result.
@@
Table 1 tests whether footprint-matched structural damage propagates to downstream document use. Rather than introducing a separate document QA benchmark, we reuse the same 975 clean-derived QA pairs under four counterfactual parser outputs, so that only the perturbation condition changes. For retrieval, the corpus is restricted to chunks from the same page and uses the same block-aware chunking across all conditions; this isolates whether answer-bearing evidence remains available after parsing. The key contrast controls for visual footprint: area-matched erasure (AM) and the structural probe (Str.) perturb nearly the same page area, with TOR around 0.10–0.11%.
@@ 5.2.3 B-SLR Channels and Faithfulness
Returning to the parser-level diagnostic, Table 2 decomposes composite B-SLR into an IoU-fail channel, where no valid geometric match survives, and a text-fail-only channel, where geometry is retained but textual identity is lost. Text-side failure is non-redundant and dominant in both pipelines: it exceeds the IoU channel on MinerU (.043 vs. .028) and more strongly on PP-V3 (.085 vs. .026), showing that many structural failures preserve coarse geometry while losing recognized content.
The channel coupling differs sharply across architectures. MinerU shows coherent geometry-text variation ($\rho_{s}=.891$), whereas PP-V3 is driven mainly by the text channel ($R^{2}_{\mathrm{text}}=.900$ vs. $R^{2}_{\mathrm{IoU}}=.186$). Thus, B-SLR does not collapse to a single text-only score: it captures architecture-specific geometry–text failure coupling while remaining aligned with OCR instability, reaching $R^{2}=.727/.916$ and Spearman $\rho=.911/.973$ on MinerU/PP-V3.
5.2.4 Exposure Descriptors Predict Failure Pathways
@@
Parser Policy Exposure / Budget Damage Failure Pathway Efficiency
@@
Table 4: Phase 2 policy audit. Under the shared probe space and page pool, targeted policies reveal the expected efficiency pattern: structural loss is dominated by topology-sensitive failures and is achieved with much smaller footprint budgets than random sampling. Metric details are provided in Appendix B.7.
@@ 5.2.5 Sensitivity Across Probe Families
The NT sweep further reveals parser-dependent sensitivity to sustained structural interference. MinerU increases gradually as the target $\tau$ increases, whereas PP-StructureV3 shows a steeper rise in topology-dominant loss at higher target levels. Overall, Phase 1 shows that the probe space contains qualitatively different failure regimes: severe erasure produces miss-heavy spikes, while structure-targeted interference increasingly induces topology-dominant failure.
@@ 5.2.6 Targeting Efficiency Across Policies
Table 4 reports the Phase 2 policy audit, where all policies share the same probe space and page pool, so differences reflect targeting logic rather than attack strength. Targeted policies consistently trade footprint for structural contact: Rule-based matches Random on PP-V3 B-SLR with about one seventh of the TOR, and VLM/Rule-based achieve the highest per-footprint efficiencies on MinerU/PP-V3. Their high TopoShare and near-zero Miss indicate that this efficiency mainly comes from topology-sensitive disruption rather than direct physical occlusion. Thus, under a shared $\mathcal{S}$, vulnerability depends not only on how much area a policy perturbs, but on whether it places probes where layout structure is sensitive; Appendix D.6 shows the same rank-consistency trend.
@@ 6 Conclusion
We studied how document parsers break under controlled visual perturbations and identified Footprint Bias: pixel footprint alone can mis-rank structural vulnerability. We proposed ProSA, a lightweight output-level framework that decouples controlled probing, policy-guided selection, and structure-aware diagnosis. Across MinerU and PP-StructureV3, B-SLR tracks OCR instability more faithfully than affected area, exposure descriptors separate occlusion- and topology-dominant pathways, and targeted probes reveal disproportionate QA/retrieval degradation under small area budgets. Overall, DLA robustness should be audited through parser-visible structural loss and failure pathways, not footprint alone.
@@
This work focuses on diagnostic auditing rather than universal robustness certification. First, our experiments instantiate the framework on representative DLA pipelines and layout benchmarks; extending the audit to broader parser families and document domains is a natural direction for future work. Second, our current implementation uses explicit layout elements and recognized text as the shared output interface; adapting the framework to less structured document understanding systems may require additional output alignment modules.
@@ Appendix A Implementation Details of Probe and Policy Spaces
Probe Geom. Behavior Key parameters Placement $\mathcal{T}$
P1 (horiz. crease) line inject $w\!\in\![1,10],\;l_{r}\!\in\![0.5,1.0]$ a/c/r
P2 (vert. crease) line inject $w\!\in\![1,10],\;l_{r}\!\in\![0.5,1.0]$ a/c/r
P4 (rect. modification) rect erase $a_{\mathrm{area}}\!\in\![3\%,25\%],\;\beta\!\in\![0.2,1.0]$ c/b
P5 (thin horiz. line) line inject $w\!\in\![1,5],\;l_{r}\!\in\![0.2,0.8]$ b/c/r
P8 (irregular patch) blob blend $r_{b}\!\in\![30,80],\;\kappa\!\in\![0.1,0.5],\;\alpha\!\in\![0.3,0.7]$ a
@@ A.2 Probe Generation and Placement
This subsection specifies how the probe catalog in Table 5 is rendered on a document page. We denote geometry masks by $\mathcal{H}_{(\cdot)}$ to distinguish probe geometry from the diagnostic exposure families in the main text. The geometry primitives are implemented as binary masks:
$\mathcal{H}_{\mathrm{line}}(\theta,l,w)$ $=\{(x,y):$
$\phantom{={}}|x\sin\theta-y\cos\theta|<\tfrac{w}{2},$
$\phantom{={}}0\leq x\cos\theta+y\sin\theta\leq l\},$ (20)
@@
$\mathcal{H}_{\mathrm{rect}}(w_{r},h_{r})$ $=\{(x,y):|x|\leq w_{r}/2,$
$\phantom{={}}|y|\leq h_{r}/2\},$ (22)
$\mathcal{H}_{\mathrm{blob}}(r_{b},\kappa)$ $=\{(r,\phi):r\leq r_{b}\bigl(1+\kappa\cdot$
$\phantom{={}}\mathrm{Perlin}(\phi)\bigr)\},$ (23)
$\mathcal{H}_{\mathrm{point}}(n,r,\sigma)$ $=\bigcup_{i=1}^{n}\mathcal{H}_{\mathrm{disk}}(r;c_{i}),$
$\phantom{={}}c_{i}\sim\mathcal{N}(\mu,\sigma^{2}I).$ (24)
@@
$I^{\prime}(x,y)=\bigl(1-\alpha(x,y)\bigr)I(x,y)+\alpha(x,y)q(x,y),$ (25)
where $q(x,y)$ is the probe texture or background target. The three composition behaviors in Table 5 are instantiated as inject ($\alpha=1$), blend ($\alpha\in(0,1)$), and erase (blending toward the local background color).
@@
$M_{\mathrm{anchor}}=\operatorname{Dil}\!\left(\bigcup_{i}\partial b_{i},\delta\right)\setminus M_{\mathrm{content}}^{\mathrm{erode}(\delta)},$ (26)
@@ A.3 Policy Families
$\pi(\text{context})=\hat{P}=\langle\hat{\mathcal{H}},\hat{\mathcal{V}},\hat{\mathcal{B}},\hat{\mathcal{T}}\rangle$, (27)
where the context features are extracted from the current page, and $\hat{P}$ follows the same geometry–appearance–behavior–placement schema as Section A.1.
@@ B.1 Exposure Descriptor Definitions
This subsection gives the operational definitions of the five exposure descriptors used in Section 4.2. Let $A(P)$ denote the perturbation mask induced by probe configuration $P$, and let $(W,H)$ be the page size. For annotated layout elements $L$, we define the annotated layout support and boundary support as
@@
$M_{\partial L}$ $=\operatorname{Dil}\!\left(\cup_{\ell\in L}\partial b(\ell),\delta\right)\setminus\Omega_{L}^{\mathrm{erode}(\delta)}.$
@@
$\mathrm{TOR}(P)=\frac{|A(P)|}{W\times H},$
$\mathrm{ACR}(P,L)=\frac{|A(P)\cap\Omega_{L}|}{|\Omega_{L}|},$
$\mathrm{BPO}(P,L)=\frac{|A(P)\cap M_{\partial L}|}{|M_{\partial L}|},$
$\mathrm{BOC}(P,L)=\frac{|\{\ell\in L:b(\ell)\cap A(P)\neq\emptyset\}|}{|L|},$
$\mathrm{EIR}(P,E)=\frac{|\{e\in E:b(e)\cap A(P)\neq\emptyset\}|}{|E|}.$ (29)
@@ B.2 Output Matching and TextSim
$\mathrm{IoU}(e,m(e))$ $\geq\tau_{\mathrm{iou}},$ (31)
$\mathrm{TextSim}(e,m(e))$ $\geq\tau_{\mathrm{text}}.$
@@
Before computing text consistency, both strings are stripped of leading/trailing whitespace and case-folded to lowercase; internal whitespace is preserved, and no punctuation removal or tokenization is applied. Let $\tilde{a}$ and $\tilde{b}$ denote the normalized strings. We compute text consistency by character-level longest common subsequence:
$\mathrm{TextSim}(a,b)=\begin{cases}1,&|\tilde{a}|=|\tilde{b}|=0,\\[2.0pt] \frac{|\mathrm{LCS}(\tilde{a},\tilde{b})|}{\max(|\tilde{a}|,|\tilde{b}|)},&\text{otherwise}.\end{cases}$ (32)
TextSim is used only as a matching criterion for B-SLR, not as the CER computation itself.
@@ B.4 Terminal Metric Definitions
For the OCR channel, we use character error rate (CER) as the primary terminal judge. Because the audit compares clean and perturbed parser outputs, the clean parser text serves as the reference text for each matched element. Using the maximum-IoU correspondence defined in Appendix B.2, let $r_{e}=s(e)$ and $h_{e}=s(m(e))$. The element-wise CER is
@@
$E_{\mathrm{text}}=\left\{e\in E\;\middle|\;s(e)\neq\epsilon\right\}$ (35)
@@
$\Delta_{\mathrm{det}}$ $=\Delta\mathrm{mAP}$ (38)
$=\mathrm{mAP}(\mathcal{Y},E)-\mathrm{mAP}(\mathcal{Y},E_{\mathrm{adv}}).$
@@ B.5 Operational Constants and Threshold Rationale
$\rho(e,P)=\frac{|b(e)\cap A(P)|}{|b(e)|}$ (39)
measures the fraction of an element box directly covered by the perturbation mask. A structurally failed element with $\rho(e,P)\geq\eta_{\mathrm{occ}}$ is assigned to the direct-occlusion subset $\mathcal{U}_{\mathrm{miss}}$; otherwise, it is assigned to the topology-dominant subset $\mathcal{U}_{\mathrm{topo}}$.